OpenStudy (karuna):

2x^2+8x+3 how do i find the vertex of this function

OpenStudy (jdoe0001):

you can always find the vertex of a parabola function like \(\bf y = ax^2+bx+c\) at the coordinates of \(\bf \left(-\cfrac{b}{2a}\quad ,\quad c-\cfrac{b^2}{4a}\right)\)

OpenStudy (karuna):

i can use any formula for the function

OpenStudy (jdoe0001):

yes

OpenStudy (jdoe0001):

well, you can use any "coefficients" for that, yes

OpenStudy (karuna):

so would the answer of this function be 2/1

OpenStudy (anonymous):

Hello Everybody, I am in 3rd grade and for science there is a question asking me how parents make a baby? I dont know plz help. i will give a medal

OpenStudy (jdoe0001):

well... dunno.. what did you get for \(\bf \left(-\cfrac{b}{2a}\quad ,\quad c-\cfrac{b^2}{4a}\right)\quad ? \)

OpenStudy (anonymous):

do you have a graphing calculator? if so graph it and the lowest point on the graph will be the vertex

OpenStudy (anonymous):

plz hellp

OpenStudy (karuna):

When the mans sperm reaches the egg

OpenStudy (karuna):

How do i find the y-intercept of the graph in a function

OpenStudy (karuna):

so would the answer of this function be 2/1

OpenStudy (karuna):

i can use any formula for the function

OpenStudy (jdoe0001):

you can always find the y-intercept by setting x = 0 and solving for "y"